Run Meeting: August 21st, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings

12:00PM RC Meeting (https://bluejeans.com/7576835804)

RC: M. Hattawy

2020

March
19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30,31
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21 22 23 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21


Summary of Aug. 21th

Owl shift

  1. Shift summary by the shift expert: shift was cancelled.
  2. Issues: NL vacuum issue.
  3. Recorded data: 0

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3831852
  2. Issues: NO beam because of the NL vacuum issue.
  3. Recorded data: zero

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3831991 . Reestablished beam on tagger magnet. Harp scans https://logbooks.jlab.org/entry/3832026 and https://logbooks.jlab.org/entry/3832058 . Took 2 Moller runs with HWP in, then out. Degaussed tagger magnet and brought beam on target.
  2. Issues: Moller quads power supply tripped (too hot?) - ended Moller run. We were able to restore the PS after a while.
  3. Recorded data: Polarization (%) 82.488 +/- 1.49 for HWP IN, -84.400 +/- 1.652 for HWP OUT.

Whole day summary

  1. Total recorded data today: 0
  2. Part-II total accumulated data on D2: 754M
  3. Concerns: none


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded